Exercise 15.76 Part A A mixture of water and graphite is heated to 600 K. When the system comes to equilibrium it contains Find the amount of CO in the flask when the system returns to equilibrium. Express your answer to two significant figures and include the appropriate units. 0.12 mol of H2, 0.12 mol of CO, 0.37 mol of H20, and some graphite. Some 02 is added to the system and a spark is applied so that the H2 reacts completely with the O2. ncoValue Units Submit My Answers Give Up
Answer #1

C(graphite) + H2O(g) ---> H2(g) + CO(g)

K = nCO*nH2/nH2O

    = 0.12*0.12/0.37

   K = 0.039

0.039 = (0.12+x)*x/(0.37+0.12-x)

x = no of mol of CO when equilibrium returns = 0.080 mol